OpenStudy (anonymous):

How does π^(ie)+1=0 work? Thnx

OpenStudy (anonymous):

Did you mean \[e^(i \pi)+1=0\]?

OpenStudy (anonymous):

yes, sry, I mixed it up :S

OpenStudy (anonymous):

Euler gives us the identity: \[e^(ix) = \cos(x) + \sin(x)\] so when you sub pi in as the radian angle it become an evaluation of sin and cos at 180 degrees \[\cos(\pi) + \sin(\pi)\ + 1 = 0\] [-1] + [0] + 1 = 0 which is true.

OpenStudy (anonymous):

thanks, that makes sense :)